Circular No. 7362

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

POSSIBLE NOVA IN SAGITTARIUS S. Nakano, Sumoto, Japan, reports the discovery of a possible nova

(mag 10.5) by Yukio Sakurai, Otsuka-cho, Mito, on two Fuji Color Super G400 films taken with a 300-mm f/2.8 lens on Feb. 4.863 UT. A prediscovery film by Sakurai on Jan. 25.863 shows the star at mag 11.5. It was not present on his films taken between 1991 and 2000 Jan. 20 (limiting mag 10.8 on Jan. 20). A CCD frame taken with a 0.40-m f/6 reflector on Feb. 6.812 by Y. and R. Kushida, Yatsugatake South Base Observatory, shows the star at mag 10.4 and yields the following position: α = 17^{h}55^{m}09^{s}.84, δ = −19^{o}46′01′′.0 (equinox 2000.0). J. Mattei, AAVSO, reports that no object was found by K. Malatesta, AAVSO, at the position of the new star on the Digital Sky Survey (SERC-V Equatorial Extension, epoch 1987.336).

SUPERNOVA 2000I IN NGC 2958 T. Puckett, Mountain Town, GA; and D. George, Ottawa, ON, report

the discovery of an apparent supernova (mag 18.6) on an unfiltered CCD frame (limiting mag 20.3) taken with the Puckett Observatory 0.60-m automated supernova patrol telescope on Feb. 4.17 UT. SN 2000I is located at α = 9^{h}40^{m}41^{s}.54, δ = +11^{o}53′08′′.5 (equinox 2000.0), which is 1′′.8 west and 10′′.2 south of the center of NGC 2958. The new object was also present on an unfiltered CCD frame taken on Feb. 6.34, but it does not appear on Palomar Sky Survey images taken on 1995 Mar. 29 (limiting mag ∼ 21.0) and on 1954 Feb. 26 (limiting mag ∼ 20.1). A. Rest and K. Krisciunas, University of Washington, report an observation of SN 2000I on Feb. 10.32 with the Apache Point Observatory 3.5-m telescope, yielding V = 17.85 ± 0.06, B − V = +0.84 ± 0.04.

SUPERNOVA 2000J IN UGC 8510 Puckett and George also report the discovery of an apparent super-

nova (mag 19.5) in UGC 8510, found on an unfiltered CCD frame (limiting mag 20.5) taken as above on Feb. 4.37 UT. SN 2000J is located at α = 13^{h}31^{m}16^{s}.24, δ = +29^{o}22′32′′.3 (equinox 2000.0), which is 16′′.9 west and 28′′.3 north of the center of UGC 8510. The new object was also present on an unfiltered CCD frame taken on Feb. 6.24, but it does not appear on Palomar Sky Survey images taken on 1989 Mar. 5 (limiting mag ∼ 21.1) and on 1950 June 5 (limiting mag ∼ 20.0). Rest and Krisciunas report that an observation of SN 2000J made (as above) on Feb. 10.30 yields V = 18.52 ± 0.10, B − V = +1.14 ± 0.08.

2000 February 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green